NvidiaIntely (nvidiaintely):

Find the height of a cylinder with radius 25 cm and surface area 8953.54 cm2. 32.00 cm 16.00 cm 31.00 cm 62.00 cm

OpenStudy (danjs):

Recall , the surface area for a cylinder is 2 circles for the top and bottom, and the side wall added up.. The circles area - pi*r^2 The side area is circumference times height - 2*pi*r * h

OpenStudy (danjs):

\[SA=2\pi*r^2+2\pi*r*h\]

OpenStudy (danjs):

right?

NvidiaIntely (nvidiaintely):

@Kainui

OpenStudy (danjs):

THey gave you surface area and the radius. SA =8953.54 r = 25 You can find the height from those... \[\large SA=2\pi*r^2+2\pi*r*h\] \[\large 8953.54 = 2\pi*(25)^2+2\pi*25*h\]